### v2

# Genesis of Axion Dark Matter via Black Hole and Cosmic String Dynamics

Abstract
The global $U_{PQ}(1)$ (Peccei-Quinn) symmetry, introduced to resolve the strong CP problem, predicts the existence of the axion—a pseudo-Nambu-Goldstone boson that is also a leading candidate for dark matter. The spontaneous breaking of this symmetry produces global strings, which decay by radiating massive axions and gravitational waves.In this study, we investigate the decay of cosmic axion strings around a Schwarzschild black hole, estimating the energy radiated due to string contraction as well as the decay timescales of the string loops. For primordial black holes (PBHs) with masses as small as $10^{-16}M_{\odot}$, the total radiated energy by the string is found to be on the order of $10^{27}$ GeV, encompassing both axion emission and gravitational waves. A key finding is that the presence of a central black hole significantly accelerates the decay of cosmic string loops, substantially reducing their lifetimes. We present these results as an initial assessment of whether axion radiation from PBH–cosmic string systems could contribute to the dark matter content in galactic halos and highlight the decay time as an important observational signature. Further detailed modeling is required to fully clarify the role of this mechanism in cosmic dark matter.
keywords:
Cosmic Strings, Axions, Peccei-Quinn Symmetry, Black holes, Dark Matter, Gravitational Waves
1 Introduction
Peccei-Quinn symmetry has been proposed in the Standard Model as a potential solution to the strong CP problem [1]. The pseudo-goldstone boson of this model, called the axion, [2] also serves as a dark matter candidate [3, 4, 5, 6, 7]. This $U_{PQ}(1)$ symmetry when spontaneously broken, leads to the formation of global strings called cosmic axion strings [8], [9]. Multiple studies have shown that global cosmic strings decay into radiation which can be massless goldstone bosons, and massive particles like axions [10], [11] and even gravitational waves (GW) [12, 13, 14].
Research on cosmic strings also involves their interaction with black holes, which lead to interesting effects. Strings when attached to a rotating black hole have been proposed to extract the black hole’s rotational energy [15], [16]. Further, string interactions with primordial black holes (PBHs) suggest the formation of a large black hole-cosmic string network [17]. Depending on their trajectories, strings near black holes may be chaotically captured or scattered [18], [19], while circular strings can exhibit vibrational and oscillatory behavior around a Schwarzschild black hole [20], [21]. Further, recent studies have derived the modified relativistic orbits for cosmic string - black hole systems [22]. Previous studies by the authors have also explored cosmic string–black hole systems in low-mass X-ray binaries, suggesting that the influence of strings on black hole spin and the system’s orbital period and accretion structure could produce observable signatures, providing potential evidence for the existence of cosmic strings [23, 24, 25]. Recent work has also examined the role of primordial black holes (PBHs) as a potential source of axionlike particles (ALPs) through Hawking radiation [26].
In this manuscript, we study the effect of a black hole on a circular cosmic axion string around it and estimate the energy radiated in the form of axions due to the contraction of the string. We establish the equation governing the radiation and the resultant increase in energy density, and then analyse the influence of various parameters. The energy radiated by this mechanism is found to be large, ranging from $\sim 10^{25}$ GeV up to $\sim 10^{53}$ GeV for increasing black hole mass and axion decay constant. When these values are compared with observational dark matter estimates, such as the Milky Way’s halo mass of order $\mathcal{O}(10^{68})$ GeV, the results suggest that this mechanism falls short of fully accounting for the dark matter content in galactic halos. A significant finding of this study is that the presence of a central black hole considerably accelerates the decay of cosmic string loops, thus drastically reducing their decay time and impacting the overall dynamics of axion emission. These preliminary findings motivate further investigation, particularly into the co-emission of gravitational waves as a promising observational channel. All calculations are performed assuming natural units ($\hbar=c=1$) throughout the manuscript.
2 Circular string around a black hole
The energy per unit length for a static, straight axion string is has been shown [8] as
$\mu=2\pi f_{a}^{2}\ln(\Lambda/\delta)$ (1)
where $f_{a}$ is the axion decay constant, $\delta$ is the string core radius (or thickness). The cutoff radius $\Lambda$ for a closed loop, (in our case a circular string) is the size of the loop [27], which we take
$\Lambda=2\pi r$ (2)
with $r$ being the radius of the circular string.
Figure 1: Depiction of the spherical cell considered for the cosmic string-black hole system
Now, we consider a cell of spherical shape having radius $r$ and one string loop of the same radius inside it for spatial symmetry as illustrated in Fig. 1 (Image not to scale). The red arrows indicate the direction of the contraction of the cell with the contracting string. It can be easily calculated that the energy inside the cell is
$E(r)=4\pi^{2}f_{a}^{2}r\ln(2\pi r/\delta)$ (3)
As the string loop contracts, kinks appear causing the axion emission by the string to maintain the energy inside the contracting cell [8],[9]. Hence, the energy of radiated axions evaluated (at some new radius $r^{\prime}<r$) as
$\Delta E$ $=E(r)-E(r^{\prime})$
$=4\pi^{2}f_{a}^{2}\left(r\ln(2\pi r/\delta)-r^{\prime}\ln(2\pi r^{\prime}/\delta)\right)$
$=4\pi^{2}f_{a}^{2}\Delta(r\ln(2\pi r/\delta))$ (4)
Further, a schwarzschild black hole of mass $M$ has been taken at the center of the cell such that the center of the black hole coincides with the center of the cell. The schwarzschild metric is now given by [20],
$ds^{2}=-(1-2GM/r)dt^{2}+(1-2GM/r)^{-1}dr^{2}+r^{2}(d\theta^{2}+sin^{2}\theta d\phi^{2})$ (5)
where the authors have taken (by ansatz)
$x^{0}=t=t(\tau),\quad x^{1}=r=r(\tau),\quad x^{2}=\theta=\pi/2,\quad x^{3}=\phi=\sigma.$ (6)
for the worldsheet coordinates $(\tau,\sigma)$ of the string. From (5), we get the equations of motion as
$\dot{t}=\frac{E_{m}}{1-2GM/r}$ (7)
$\dot{r}^{2}=E_{m}^{2}-r^{2}(1-2GM/r)$ (8)
where $E_{m}$ is the energy of the string at maximal radius. The maximal radius is given by the solution of these equations at $\tau=0$ as
$r_{max}=r(\tau=0)=GM+\sqrt{G^{2}M^{2}+E_{m}^{2}}$ (9)
matching with the results obtained in [20]. The contraction of the string continues while radiating axions till it reaches the event horizon, giving us the minimum radius as
$r_{min}=2GM$ (10)
Furthermore, using (7), the coordinate time taken for the string to contract from $r_{max}$ to $r_{min}$ has been calculated as
$t_{\rm BH}(\tau)=E_{m}\Delta\tau+2GM\log\left|\frac{\tan(\Delta\tau/2)+\epsilon}{\tan(\Delta\tau/2)-\epsilon}\right|$ (11)
where $\epsilon=(\sqrt{G^{2}M^{2}+E^{2}}-GM)/E_{m}$ and the world-sheet time $\Delta\tau$ is given by
$\Delta\tau=cos^{-1}\left(\frac{GM}{\sqrt{G^{2}M^{2}+E_{m}^{2}}}\right)$ (12)
Now, by substituting these extreme values of radius from (9) and (10) in (4), the energy radiated by the string during contraction has been calculated as,
$\Delta E=\;$ $4\pi^{2}f_{a}^{2}\bigg((GM+\sqrt{G^{2}M^{2}+E_{m}^{2}})\ln\left(\frac{2\pi(GM+\sqrt{G^{2}M^{2}+E_{m}^{2}})}{\delta}\right)$
$\quad-2GM\ln\left(\frac{4\pi GM}{\delta}\right)\bigg)$ (13)
and the energy emission rate can be shown to be $\Delta E/\Delta\tau$. Further, it can also be established that the energy density inside the cell increases as
$\Delta\rho(r)=\frac{\Delta E}{\frac{4}{3}\pi r^{3}}=\frac{3\pi f_{a}^{2}}{r^{3}}\Delta(r\ln(2\pi/\delta))$ (14)
Interestingly, we have found that the energy of the radiated axions and the rate of radiation remains nearly constant in this system with respect to the fixed mass of the black hole, the initial (maximum) energy and decay constant of the string. These results imply that such systems could generate a continuous background of axion radiation, potentially contributing to the dark matter density in galactic halos.
3 Numerical analysis
3.1 Energy Radiation
With the equations for the radiated energy established in the previous section, it is important to get an estimate on the values of $\Delta E$. These values will help us in understanding the importance of this mechanism and the significance of its contribution to the energy of the universe.
In this regard, we simulate $\Delta E$ over a mass range starting from $10^{-16}M_{\odot}$ of primordial black holes (PBHs) upto the largest $10^{10}M_{\odot}$ of supermassive black holes (SMBHs) for a decay constant $f_{a}=10^{10}$ GeV depicted in Fig. 2 (both axes are in logscale). Similarly, Fig. 3 further examines energy dependence on decay constant in the range $f_{a}\in(10^{9},10^{12})$ GeV [3] by taking $M=10M_{\odot}$. The string core radius corresponds to $\sim 1/f_{a}$. Simulations and plots were implemented using Python’s Matplotlib library.
Figure 2: $\Delta E$ for varying black hole mass $(M)$
Figure 3: $\Delta E$ for varying decay constant ($f_{a}$)
Figure 4: $\Delta E$ for varying black hole mass $(M)$ and decay constant $f_{a}$
Fig. 2 and Fig. 3 demonstrate the linear relationship between energy gain and mass and decay constant as expected from (13) . It is interesting to note that even for the smallest black holes, the energy radiated is as high as $\sim\mathcal{O}(10^{27})$ GeV. It is emphasized here that the order of the energy radiated is affected when a different decay constant or black hole mass is considered, but the plot’s behavior remains unchanged. Fig.4 illustrates the same behavior by displaying the combined effect of variations in the decay constant and the black hole’s mass as in (12).
Figure 5: $\Delta E$ corresponding to the GWs (green dotted line) and particle emission (purple line) for varying mass
The energy calculated and plotted in Fig. 2, Fig. 3 and Fig. 4 illustrates the total energy released by this mechanism. String loop decay primarily dissipates energy through two channels: particle emission, specifically axions in this context, and gravitational radiation. However, recent findings [28] indicate that power radiated GW ($P_{\rm GW}$) is considerably suppressed in comparison to power radiated in particle emission ($P_{\varphi}$) quantified as
$\frac{P_{\rm GW}}{P_{\varphi}}\approx\mathcal{O}(10)\left(\frac{v}{m_{\rm p}}\right)^{2}\ll 1.$ (15)
where $v$ is the vacuum expectation value associated to string formation and $m_{\rm p}$ is the Planck mass. Applying this relation (15) in our framework allows a clear separation of the energy into gravitational wave and particle radiation This distinction is depicted in Fig. 5, demonstrating the dominance of particle radiation.
3.2 Decay Time
The decay time for our black hole - cosmic string system is estimated using (11) and (12) for an observer at infinity. In the absence of black hole however, the decay time depends linearly on the loop length $L$ as described by [11]
$t_{\rm free}\sim 1.4L$ (16)
measured in string’s rest frame. Incorporating the Lorentz boost factor of 1.25, Fig. 6 compares decay times with and without the central black hole, clearly showing accelerated decay induced by the black hole’s gravitational field. An additional noteworthy observation is the higher decay time up to a loop radius of 8 $GM$, illustrated in Fig. 7, which arises from the gravitational time dilation caused by the black hole. This effect has significant implications for potential observational signatures of the system.
Figure 6: Decay time with (blue line) and without (red dashed line) a black hole for varying black hole mass
Figure 7: Decay time/loop radius with (blue line) and without (red dashed line) a black hole for varying black hole mass depicing the increased decay time close to the black hole
4 Results and discussions
In order to interpret the physical implications of the energy radiated from circular cosmic axion strings contracting around black holes, we perform a quantitative analysis of the emitted energy over a broad range of black hole masses and axion decay constants. This approach aims to investigate whether the energy output from this mechanism can account for the dark matter content observed in galactic halos. Our simulations cover a wide range of black hole masses from primordial black holes ($\sim 10^{-16}M_{\odot}$) to supermassive black holes ($\sim 10^{10}M_{\odot}$) with typical axion decay constants between $10^{9}$ and $10^{12}$ GeV.
The total radiated energy ranges from approximately $10^{27}$ GeV for the smallest black holes to as high as $10^{53}$ GeV for supermassive black holes, indicating a substantial release of axion radiation in these systems. To provide context, the estimated dark matter mass in the Milky Way halo is about $10^{11}M_{\odot}$ or $\sim 10^{68}$ GeV. Our results show that the radiated energy from these mechanisms spans several orders of magnitude below but approaches the Galactic dark matter scale, suggesting a potentially significant but not definitive contribution to the dark matter content in galactic halos.
Significantly, our analysis reveals that the decay times of cosmic string loops are considerably shortened by the gravitational influence of the central black hole. This accelerated decay serves as a notable effect that could impact both the dynamics of axion emission and possible observational signals, especially through gravitational wave co-emission. Gravitational wave emission, although suppressed relative to axion particle emission, is non-negligible and represents a promising channel for observational signatures of this mechanism.
These findings prompt further detailed modeling, particularly accounting for evolving black hole masses, decay time effects, and string loop dynamics over cosmological timescales, to fully assess whether axion radiation from these systems can quantitatively explain observed dark matter halos.
As discussed previously, it has been shown that [26] future gamma-ray telescopes such as e-ASTROGAM could detect ALPs emitted by PBHs if these particles decay to photons before reaching Earth, producing distinctive double-peaked spectral signatures compared to Standard Model predictions. This method of photon coupling could similarly be applied in detecting the axions radiated due to string contraction in black hole spacetime. Additionally, as PBHs grow into SMBHs, the parameters determining loop contraction become dynamic, further affecting the total energy radiated. These factors highlight that the scenario described does not capture the complete picture and must be refined to account for evolving nature of PBH mass and string loop dynamics.
5 Conclusion
This study investigates a Schwarzschild black hole encircled by a circular cosmic axion string, where the string’s equation of motion and extremal values for loop size have been analysed. As the string contracts, developing kinks, it emits axion radiation, with energy output governed by the black hole’s mass, loop size, and decay constant. The results quantify emitted energies across PBHs and SMBHs for a range of axion decay constants, providing a numerical basis to explore the extent to which these processes may contribute to galactic dark matter halos.
While the emitted axion energy shows promising magnitudes, it remains an open question whether this mechanism can fully account for the dark matter observed in galaxies. The presence of a central black hole significantly influences decay times and energy output, suggesting potential observational signatures in both particle and gravitational wave channels. Future work will need to incorporate more complex string configurations and evolving system parameters to rigorously test the viability of axion emission from black hole-cosmic string systems as a dominant dark matter genesis mechanism.